Consider a makeup mirror that produces a magnification of 1.3 when a person’s face is 13 cm away.
What is the focal length of the makeup mirror in meters?
1
Expert's answer
2020-07-20T15:04:41-0400

To get focal length first we need to find the image distance which can be obtained by the magnification formula

"m =-\frac{image distance} { object distance}"


substituting the values given

"1.3=-\frac{imagedistance}{0.13}"


image distance v = -0.169 meters (The negative values imply that the image is virtual)


Now that the value of v is known, the focal length can be computed by the lens formula

"\frac{1}{f}=\frac{1}{u} +\frac{1}{v}"


substituting the values n the formula


"\frac{1}{f}=\frac{1}{0.13 }+\frac{ 1}{0.169}"


"\frac{1}{f}=\frac{2300}{169}"


"f= 0.0735m"


Answer= 0.0735m